OpenStudy (anonymous):

What is the value of the expression 4/2^-2? A. ½ B. 1 C. 8 D. 16

OpenStudy (anonymous):

@Studyhelp_00002

OpenStudy (anonymous):

@GracieBugg

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

@amberosales

OpenStudy (anonymous):

plz help me

OpenStudy (tkhunny):

Show your work. Stop tagging the world.

OpenStudy (anonymous):

im trying im stuck

OpenStudy (phi):

do you know how to "flip" \[ \frac{1}{2^{-2} }\] ?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

See, you could have shown that.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

it is 4/2^-2

OpenStudy (phi):

what do you get when you flip \[ \frac{1}{2^{-2}} \]?

OpenStudy (tkhunny):

No, you wrote the problem statement. You did not show that you knew how to "flip" negative exponents. You could have shown that.

OpenStudy (anonymous):

ok it is

OpenStudy (anonymous):

i for got show me plz

OpenStudy (phi):

the "rule" is flip the fraction (bottom becomes top and vice versa) *and* change the exponent by multiplying the exponent by -1

OpenStudy (phi):

can you do that for \[ \frac{1}{2^{-2}} \]

OpenStudy (phi):

here is an example \[ \frac{1}{5^3}= \frac{5^{-3}}{1}\]

OpenStudy (anonymous):

ok

OpenStudy (phi):

do it in two steps: 1) first, can you write the "flipped" version of 1/2^(-2) ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

yes

OpenStudy (phi):

can you post step 1) flip 1/2^(-2)

OpenStudy (anonymous):

\[\frac{ 1 }{ 2^3 }\]

OpenStudy (anonymous):

-3

OpenStudy (phi):

start with \[ \frac{1}{2^{-2}} \] now make the bottom (the \( 2^{-2}\) ) the top and make the top (the 1) the new bottom what do you get ?

OpenStudy (anonymous):

it is the same

OpenStudy (phi):

if you had \[ \frac{a}{b}\] the flipped version is \[ \frac{b}{a} \]

OpenStudy (anonymous):

im sorry somthing so simple is hard for me

OpenStudy (phi):

ok, practice with this x/y what is that flipped?

OpenStudy (anonymous):

x/y-y?

OpenStudy (phi):

x/y flipped becomes y/x try c/d what is that flipped?

OpenStudy (anonymous):

oooo d/c

OpenStudy (phi):

yes. now try 3/2 flipped is ?

OpenStudy (anonymous):

2/3

OpenStudy (phi):

yes now this one (4*2)/(5*6) flip that

OpenStudy (anonymous):

2*4/6*5

OpenStudy (anonymous):

1/(2-^2)

OpenStudy (phi):

try 1/(2*3) flip that

OpenStudy (anonymous):

humm 1/(3*2)

OpenStudy (phi):

no, (2*3)/1 just like a/b becomes b/a c/d becomes d/c x/y becomes y/x (2*3)/(4*5) becomes (4*5)/(2*3) do you see the pattern ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

now try 1/(2*3)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok (2*3)/1

OpenStudy (phi):

yes. now use that same idea on 1/(2^-2) flip this one

OpenStudy (anonymous):

(2_^2)/1

OpenStudy (phi):

(2^-2)/1

OpenStudy (anonymous):

o woops

OpenStudy (phi):

now when we flip things they change 1/2 is different from 2/1 but if we have 1/(2^-2) and flip it to 2^-2/1 then *change the exponent* so that we have 2^2/1 it turns out that \[ \frac{1}{2^{-2}} = \frac{2^2}{1} \]

OpenStudy (phi):

so we do two things: flip and *change the exponent*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

humm

OpenStudy (phi):

now your problem is \[ 4 \cdot \frac{1}{2^{-2}} \] and 1/2^-2 is 2^2, so it is now \[ 4 \cdot 2^2 \]

OpenStudy (phi):

2^2 means 2*2 4* 2^2 means 4*2*2

OpenStudy (anonymous):

ok nvm i got it u r just beating around the boosh sorry i figured it out thx thow

OpenStudy (phi):

you should get 16